The demand and supply functions for a good is given as Q = 30 - 2p and Q = 20 + 2p. What will be the equilibrium quantity?


1
Expert's answer
2021-04-20T07:50:11-0400

To determine the equilibrium quantity the Quantity demanded must be equal to the Quantity supplied

Qd= As

Therefore;

30-2p= 20+2p

30-20=2p+2p

10= 4p

p=2.5

Substituting p to the the Qd

30-2(2.5)

30-5=25

Qe= 25units